# Putnam-like dataset summary: LLMs as mathematical competition contestants

### v1

@@ Abstract.
In this paper we summarize the results of the Putnam-like benchmark published by Google DeepMind. This dataset consists of 96 original problems in the spirit of the Putnam Competition and 576 solutions of LLMs. We analyse the performance of models on this set of problems to verify their ability to solve problems from mathematical contests.
@@ 1. Introduction
In recent months, general-purpose large language models have significantly progressed in mathematical reasoning. It was observed that LLMs are obtaining high scores in final-answer benchmarks (see [4]). In such benchmarks, models are asked only for the final answer to a mathematical problem, which is then compared to the correct one. Models don’t produce complete, rigorous solutions (or they are not verified), but they provide only the final answer, which is usually a number. It is clear that such benchmarks are not able to capture the real mathematical capability of the model. Hence, there is a need to analyze the behavior of LLMs on mathematical proof-like problems, in particular, where the aim of the problem is to prove or disprove some property. Such problems are quite popular in mathematical competitions.
@@
On September 23rd, 2025, Google DeepMind released the dataset, called Putnam-like (https://github.com/google-deepmind/eval_hub/tree/master/eval_hub/putnam_like), with 96 mathematical problems on the level of competitions for students like the William Lowell Putnam Mathematical Competition (https://maa.org/putnam/) or the International Mathematics Competition for University Students (https://www.imc-math.org.uk/). These problems were prepared in cooperation with Google DeepMind by the authors of this article as modifications of the problems used in competitions in previous years.
@@
The dataset consists of 8 sets of problems that mimic the style of the Putnam competition, i.e., in each of these sets there are two parts (A and B) with six problems of increasing difficulty. The data published by Google DeepMind has a reach structure. For each problem there is given a proposed solution (rubrics) divided into steps worth some points, summing up to 10. Six AI models were tested on each problem: gemini-2.5-pro-03-25, gemini-2.5-flash-04-17, o3-mini-high, o4-mini-high, r1, sonnet-3.7. The models’ solutions were verified and graded by human experts (this article’s authors) and by the gemini-2.5-pro_20250718 model. All the grades, along with the comments, are published in the dataset.
@@
The goal of this paper is to summarize the LLMs’ ability to solve the problems in natural language proposed in the dataset. These problems require different mathematical skills to be solved, and the proposed solution has to be completely, fully rigorously proven, even if some numerical answer can be provided. Note that, as described in the dataset’s readme file, the models were prompted only with the problem statement, without any requirements for the quality of solutions. However, in the grading process performed by experts, the solutions were graded as being given during the mathematical contests, especially the usage of numerical calculations was prohibited, the references for the non-standard theories were required, and all the steps of the reasoning were expected to be written.
Sample Problem: Set_1/A6 Level: 6 Category: Analysis Question (prompt): Let $f:\mathbb{R}^{2}\rightarrow\mathbb{R}$ be a continuous function and let $r>0$ be fixed. Suppose that $\int_{B(x_{0},r)}f(x)\,dx=0$ for any $x_{0}\in\mathbb{R}^{2}$. Does it imply that $f\equiv 0$? Answer (generated by the model): Assume, to the contrary, that $f$ is not identically zero. Then there exists Yes, the given condition implies that $f\equiv 0$. Grade: 0 points The solution is wrong (the model shows that $f$ has to be zero, which is not true). The model assumes in addition that $f\geq 0$.
@@
This study is another validation of the LLMs on Putnam-style problems. The study of LLMs solutions to the Putnam Competition problems given in natural language was done in the recent Putnam-AXIOM benchmark [4]. There is also a very interesting study of the AI models solving reformulations of Putnam Competition problems into formal languages like Lean, Isabelle, or Coq [8]. The behaviour of LLMs was also studied on problems from other mathematical competitions, including the 2025 USA Mathematical Olympiad [7], the International Mathematical Olympiad [6] (also formalized in Lean [5]), as well as original, challenging problems [3].
@@ 2.1. Total distribution of grades
The problems from the dataset were mostly correctly solved by the models. It agrees with the other benchmarks, which show that the actual models are very good at solving problems from competitions for students. For example, in MathArena [2], the models got more than 90% of points on the problems from IMC 2025. Note that the AI efficiency on IMO 2025 problems is much lower – at most 38.1%. This is due to the different characteristics of competitions for school and university students. IMO tasks require more originality, while student competition tasks are based on creative proficiency in academic topics.
@@
In this paper we study the grades given for the natural language solutions prepared by LLMs, so it is a significantly different class of solutions than the human ones. The only comparable dataset is the Putnam-AXIOM benchmark [4], where the results are low and the best one was obtained by o1-preview model with 41.94% correct answers. However, there is a crucial difference between Putnam-AXIOM and our study. They focus on the correct boxed answer to binary verify if the problem is solved or not, while in Putam-like dataset the whole reasonings were verified and partial grades could be awarded.
To obtain a reliable comparison of the difficulty of the tasks in the dataset, the authors evaluated the solutions of the same six models on tasks from the 2024 Putnam competition in a double review. The results by model are presented in Table 2 together with average scores on the dataset.
@@
Table 2 can be considered as a reference point to describe the general difficulty of the Putnam-like dataset. Based on this premise, we conclude that it is rather easier than the actual Putnam competition. Note also that in the dataset the grades from 3 to 7 are more frequent than in the Putnam competition. It was forced by the usage of rubrics given as steps of the original solution that were needed for the automatic grading. The experts were bound by these guidelines, and for this reason, more partial points were given.
@@ 2.2. Results by problem level
The boxplot is a very popular way to visualize a distribution of a quantity. The box is drawn from the first quartile to the third quartile with a horizontal line drawn inside it to denote the median. The length of the box is called the interquartile range (IQR). The whiskers are drawn at most 1.5IQR down, starting from the first quartile, and at most 1.5IQR up, starting from the third quartile. The points out of whiskers are called outliers and are considered as untypical for considered set of values.
@@
As can be seen in Figure 3, in overview, the problems are ordered in increasing difficulty, and the problems at levels 5-6 were significantly harder for LLMs than the rest. However, problems level 4 are easier for models than problems level 3. Welch’s two sample $t$-test confirms that this difference is statistically significant ($p$-value of one-sided test is equal to 0.02).
@@ 2.3. Results by category
As we see in Figure 4, the median is lower than 10 in three categories: abstract algebra, analysis, and polynomials. The explanation of this fact requires a detailed study. We suspect that, in the case of mathematical analysis, some of the models didn’t get the maximum score because of the usage of numerical computations and unjustified approximations, although the final answer was correct. The plot also shows that problems in linear algebra were easier for models than the rest of the problems.
@@ 2.4. Statistics of results by model
Significantly, the best grades were scored by gemini-2.5-pro-03-25, and the efficiency of its faster cousin gemini-2.5-flash-04-17 is also very good, see Figure 5. On the other hand, the r1 model got the lowest grades. Brief analysis of the dataset shows that the r1 model very often provides only a sketch of a correct solution, without providing any details or justifications. Such a behavior is possible due to the lack of any additional prompt explaining that the solutions should contain all the details and justifications.
@@
The results of sonnet-3.7 are similar to those of o3-mini-high and o4-mini-high. It is worth to remark, that solutions of this model are well edited and easy to read.
@@ 3. Conclusions
Using a Putnam-like dataset, we effectively simulated a mathematics competition modeled on the William Lowell Putnam Mathematical Competition, with large language models (LLMs) as participants. The difficulty level of the problems was slightly lower than in last year’s edition of the contest.

### v2

@@ Abstract.
In this paper we summarize the results of the Putnam-like benchmark published by Google DeepMind. This dataset consists of 96 original problems in the spirit of the Putnam Competition and 576 solutions generated by LLMs. We analyze the performance of models on this set of problems to verify their ability to solve problems from mathematical contests. We find that top models, particularly Gemini 2.5 Pro, achieve high scores, demonstrating strong mathematical reasoning capabilities, although their performance was notably lower on problems from the 2024 Putnam competition. The analysis highlights distinct behavioral patterns among models, including bimodal scoring distributions and challenges in providing fully rigorous justifications.
@@ 1. Introduction
General-purpose large language models have recently shown significant progress in mathematical reasoning. It was observed that LLMs are obtaining high scores in final-answer benchmarks (see [9]). In such benchmarks, models are asked only for the final answer to a mathematical problem, which is then compared to the correct one. Models do not produce complete, rigorous solutions (or they are not verified), but they provide only the final answer, which is usually a number. It is clear that such benchmarks are not able to capture the real mathematical capability of the model, except in the case of extremely challenging problems, see [8].
Hence, there is a need to analyze the behavior of LLMs on mathematical proof-like problems, in particular, where the aim of the problem is to prove or disprove some property. Such problems are quite popular in mathematical competitions. The behavior of LLMs on proof-based problems was studied on problems from a number of mathematical competitions, including the 2025 USA Mathematical Olympiad [4, 13], the International Mathematical Olympiad [4, 12] (also formalized in Lean [11]), and the International Mathematics Competition for University Students (IMC) [4, 10]. See also [6, 7] for a wide benchmarks of LLM-generated mathematical proofs for the problems originated in mathematical competitions. The parallel topic is the benchmarking of LMMs on research-level problems that requires a rigorous proof, see [14].
@@
On September 23, 2025, Google DeepMind released the dataset, called Putnam-like (https://github.com/google-deepmind/eval_hub/tree/master/eval_hub/putnam_like), with 96 mathematical problems on the level of competitions for students like the William Lowell Putnam Mathematical Competition (https://maa.org/putnam/) or IMC (https://www.imc-math.org.uk/). These problems were prepared in cooperation with Google DeepMind by the authors of this article.
@@
The dataset consists of 8 sets of problems that mimic the style of the Putnam competition, i.e., in each of these sets there are two parts (A and B) with six problems of increasing difficulty. The data published by Google DeepMind has a rich structure. For each problem there is given a proposed solution (rubrics) divided into steps worth some points, summing up to 10. Six AI models were tested on each problem, including Google’s Gemini family (gemini-2.5-pro-03-25 and gemini-2.5-flash-04-17), models of OpenAI (o3-mini-high and o4-mini-high), Anthropic’s Claude Sonnet (sonnet-3.7) and the first generation of DeepSeek model (r1). The models’ solutions were verified and graded by human experts (this article’s authors) and by the Google DeepMind’s gemini-2.5-pro_20250718 model. All the grades, along with the comments, are published in the dataset.
@@
In this paper we summarize the LLMs’ ability to solve the problems in natural language proposed in the dataset. These problems require different mathematical skills to be solved, and the proposed solution has to be completely, fully rigorously proven, even if some numerical answer can be provided. Note that, as described in the dataset’s readme file, the models were prompted only with the problem statement, without any requirements for the quality of solutions. However, in the grading process performed by experts, the solutions were graded as being given during the mathematical contests, especially: the usage of a very precise numerical approximations was prohibited (as impossible to be done during a competition), the references for the non-standard theories were required, and all steps of the reasoning were expected to be written.
@@
In this article, we focus only on the grades given by human experts, treating them as the most accurate. The correlation of the human and automatic grades needs to be studied in future work.
This study is another validation of the LLMs on Putnam-style problems. The study of LLMs solutions to the Putnam Competition problems given in natural language was done in the recent Putnam-AXIOM benchmark [9]. Note that this dataset collects problems from the actual Putnam competition, possibly with minor symbolic modifications, so most of these problems were contained in the LLMs’ training data. The Putnam-like dataset is not affected by this type of data contamination.
There is also a very interesting study of the AI models solving reformulations of Putnam Competition problems into formal languages like Lean, Isabelle, or Coq [15].
The goal of this paper is to investigate and compare the behaviors of different language models on Putnam-like problems to characterize their distinct problem-solving styles. We show that the models differ not only in overall performance but also in how they handle problems from various mathematical categories, in the distribution of scores they achieve, and in whether they provide complete, rigorous solutions or merely sketches. The Gemini models demonstrated the most mature mathematical behavior, frequently producing detailed and fully justified proofs, with few intermediate or partially correct responses. In contrast, the OpenAI’s models exhibited more exploratory reasoning: they tended to identify key ideas and partial arguments but frequently failed to develop them into complete, formally correct proofs. The Claude Sonnet model showed a similar profile, characterized by coherent and readable solutions that occasionally contained conceptual inaccuracies or omitted critical steps. The DeepSeek’s model, on the other hand, displayed the weakest performance, typically offering sketches or outlines without sufficient justification or mathematical depth.
Our analysis also allowed us to assess the difficulty of the Putnam-like dataset itself. Based on a comparison of the models’ performance on this dataset with their results on authentic problems from the 2024 Putnam competition, we find that the difficulty level of the prepared benchmark is slightly lower than that of the actual competition.
@@ 2.1. Total distribution of grades
Most problems from the dataset were solved correctly by the models. This result is consistent with findings from other benchmarks, which show that the actual models are very good at solving problems from competitions for students. For example, in MathArena [4], the models achieved scores exceeding 90% on the problems from IMC 2025. Note that the performance of the models analyzed in MathArena on IMO 2025 problems is much lower – at most 38.1%. This is due to the different characteristics of competitions for school and university students. Student competition tasks are mostly based on creative proficiency in academic topics that are well mastered by LLMs, while the IMO problems often require more originality than advanced methods. However, despite the high difficulty of the IMO 2025 problems, state-of-the-art models demonstrated strong performance within the prescribed time limits. The advanced release of Google’s Gemini Deep Think [1], Harmonic’s Aristotle [2], and an experimental model developed by OpenAI each attained gold-medal performance, successfully solving 5 out of 6 problems. In comparison, ByteDance’s Seed-Prover [5] reached the silver-medal threshold with solutions to 4 problems.
@@
Note also that in the dataset the grades from 3 to 7 are more frequent than in the Putnam competition. This distribution was a consequence of using rubrics given as steps of the original solution that were needed for the automatic grading. The experts were bound by these guidelines, and for this reason, more partial points than in the actual Putnam competition were expected.
In this paper we study the grades given for the natural language solutions prepared by LLMs, so it is a significantly different class of solutions than the human ones. The only comparable dataset is the Putnam-AXIOM benchmark [9], where the results are low and the best one was obtained by o1-preview model with 41.94% correct answers. However, there is a crucial difference between Putnam-AXIOM and our study. They focus on the correct boxed answer to binary verify if the problem is solved or not, while in the Putam-like dataset the entire reasoning was verified and partial grades could be awarded.
To obtain a reliable comparison of the difficulty of the tasks in the dataset, the authors evaluated the solutions of the same six models on tasks from the 2024 Putnam competition, using the same grading criteria (we do not publish the model’s answers and our reviews). Note that we did not know the grading schemes used in the actual Putnam competition, so this grades cannot be directly compared with Putnam results.
The results by model are presented in Table 2 together with average scores on the dataset.
@@
Table 2 can be considered as a reference point to describe the general difficulty of the Putnam-like dataset. Based on this premise, we conclude that it is rather easier than the actual Putnam competition.
@@ 2.2. Results by problem level
The distributions of the grades we visualize using well-known boxplots and dotplots (if the sample is small).
@@
As can be seen in Figure 3, in overview, the problems are ordered in increasing difficulty, and the problems at levels 5-6 were significantly harder for LLMs than the rest. However, level 4 problems are easier for models than level 3 problems. Welch’s two sample $t$-test confirms that this difference is statistically significant ($p$-value of one-sided test is equal to 0.02).
This anomaly may be attributed to human-perceived problem difficulty. The same thing one can observe in the results from the actual Putnam competition. In fact we should consider the problems as easy (level 1 and level 2), medium (3 and 4) and hard (5 and 6) where the order in each group is not important. This also highlights a potential discrepancy between human-perceived difficulty and the actual challenges faced by LLMs, suggesting that models may excel at problems requiring extensive calculations or knowledge recall, which humans might rate as more difficult.
@@ 2.3. Results by category
As we see in Figure 4, the median is lower than 10 in three categories: abstract algebra, analysis, and polynomials. The explanation of this fact requires a detailed study. We suspect that, in the case of mathematical analysis, some of the models did not get the maximum score because of the use of numerical computations and unjustified approximations, although the final answer was correct. The plot also shows that problems in linear algebra were easier for models than the rest of the problems.
@@ 2.4. Statistics of results by model
Significantly, the best grades were scored by gemini-2.5-pro-03-25, and the efficiency of its faster counterpart gemini-2.5-flash-04-17 is also very good, see Figure 5. On the other hand, the r1 model got the lowest grades. A brief analysis of the dataset reveals that the r1 model frequently provides only a sketch of a solution, without providing any details or justifications. Such a behavior is possible due to the lack of any additional prompt explaining that the solutions should contain all the details and justifications.
@@
The results of sonnet-3.7 are similar to those of o3-mini-high and o4-mini-high. It is worth remarking, that solutions of this model are well-edited and easy to read.
@@ 3. Conclusions
Using the Putnam-like dataset, we effectively simulated a mathematics competition modeled on the William Lowell Putnam Mathematical Competition, with large language models (LLMs) as participants. The difficulty level of the problems was slightly lower than in last year’s edition of the contest.
@@
Future work will focus on a deeper analysis of the details of the models’ solutions and the errors committed. Another important aspect is the correlation between human and automatic grading, which could lead to the development of more reliable evaluation systems, as well as exploring the impact of advanced prompting techniques on the models’ ability to generate rigorous mathematical proof.